Suppose A and B are 3 × 3 matrices. Show that If B is obtained from A by adding 2 times the first row of A to the last row of A, then

det (A) = det (B)


1
Expert's answer
2020-08-25T10:40:24-0400

A: "\begin{bmatrix}
 a & b & c\\
 d & e & f\\
 g & h & i
\end{bmatrix}"


B: "\begin{bmatrix}
 a & b & c\\
 d & e & f\\
 g+2a & h+2b & i+2c
\end{bmatrix}"


"det(A)=det(B)"


"a*det(\begin{bmatrix}
 e & f \\
 h & i
\end{bmatrix})-b*det(\begin{bmatrix}
 d & f \\
 g & i
\end{bmatrix})+c*det(\begin{bmatrix}
 d & e \\
 g & h
\end{bmatrix})=a*det(\begin{bmatrix}
 e & f \\
 h+2b & i+2c
\end{bmatrix})-b*det(\begin{bmatrix}
 d & f \\
 g+2a & i+2c
\end{bmatrix})+c*det(\begin{bmatrix}
 d & e \\
 g+2a & h+2b
\end{bmatrix})"

"a(ei-fh)-b(di-fg)+c(dh-eg)=a(e(i+2c)-f(h+2b))-b(d(i+2c)-f(g+2a))+c(d(h+2b)-e(g+2a))"


"aei-afh-bdi+bfg+cdh-ecg=bfg-afh+eai-bdi-ecg+cdh"


"0=0"


Proved.